# Mavors: Multi-granularity Video Representation for Multimodal Large Language Model

### v1

@@ 1 Introduction
Long-context video modeling stands as one of the most crucial capabilities within MLLMs [6, 116, 47, 67]. This capability empowers MLLMs to proficiently manage hours-long movies, documentaries, and online video streams, all of which demand sophisticated long video processing. Recent advances in MLLMs perform well in short video understanding. However, it remains challenging to build MLLMs for processing extremely long videos (lasting for hours or even longer). The difficulty lies in how to enable MLLMs to efficiently understand the extremely long video context brought by long videos.
As shown in Figure 1, we have compared three mainstream types of video MLLMs with our method, and provided the video caption results of different methods for better illustration. Specifically, in Figure 1(a), these methods (e.g., LLaVA-Video [124], InternVL 2.5 [14]) usually employ the sparse sampling strategy to decrease the number of frames and reduce the computation costs. However, these methods have a significant limitation, where many temporal contexts are lost as many frames are not sampled. Thus, the performance results of video-related tasks, which require detailed temporal contexts from many frames, are degraded a lot for these methods. When compared to methods in Figure 1(a), some methods (e.g., Oryx [60], Qwen2VL [98]) have introduced the strategy of dense sampling with low-resolution input in Figure 1(b). However, for these methods, many spatial contexts are lost as only the low-resolution frames are given, which also significantly degrade the results of video-related tasks requiring detailed spatial contexts, e.g., video captioning. Recently, in Figure 1(c), several works (e.g., VideoLLaMA 3 [116], VideoChat-Flash [47]) have proposed token compression strategies (e.g., token merge or token dropping), which reduces tokens based on vector or pixel similarity and effectively preserves spatial-temporal features of large visual elements. However, token compression inevitably leads to the loss of information regarding small spatial objects, subtle temporal motions, and interactions among multiple objects, thereby posing challenges for understanding complex scenes.
@@ 2.1 MLLM Architecture
Current MLLMs employ two architectural strategies for visual processing. The first paradigm is based on cross-attention approach, which maintains frozen model parameters while establishing dynamic visual-language interactions through attention mechanisms [2] . Alternatively, the second paradigm processes visual content through pretrained encoders (CLIP [76], SigLIP [115]) before concatenating image tokens with text embeddings for unified language model processing [51, 54, 53, 43, 55]. The second paradigm can be readily extensible to video analysis through sequential frame processing [45, 116], and many architectural innovations for temporal modeling have been proposed [34, 103, 56].
@@ 2.2 MLLM for Video Understanding
Existing MLLMs have revealed divergent capabilities in temporal comprehension across different video durations. While existing systems demonstrate proficiency in minute-scale video analysis [45, 47, 50], emerging efforts targeting hour-level sequences [23, 101] face fundamental challenges. To address the challenges of long video modeling, current approaches primarily pursue two optimization directions: (1) context window expansion for large language models [23, 101, 120, 108] and (2) efficient token compression via spatial-temporal feature distillation [20, 49, 85, 86, 90, 104]. For the first strategy, though theoretically enabling long-sequence processing, suffers from impractical computational overhead, which bring significant challenges for practical applications. In contrast, recent token compression methods like LLaMA-VID [49] achieve compression rates at the cost of discarding subtle details, which results in performance degradation on standard video understanding benchmarks. When compared to the existing works, our Mavors can directly process the raw videos to maintain spatial and temporal details well with acceptable computation costs.
@@ 3.1 Preliminaries
Necessity of Dense Sampling with High Resolution. As shown in Figure 2 and Figure 3, we have compared the results of two popular video MLLMs (i.e., Qwen2.5-VL-7B [4] and Oryx-1.5-7B [60]) on two representative benchmarks (i.e., Video-MME [22] and DREAM-1K [96]). Specifically, the Video-MME focuses on multiple-choice question answering based on video content and requires a better understanding of the temporal relations between different frames. DREAM-1K involves open-ended video captioning, where models must generate detailed descriptions of the main events in the video. Thus, both the spatial and temporal fine-grained details are important. In Figure 2, we observe that performance increases a lot when increasing the number of frames, which shows the necessity of dense sampling with more frames. In Figure 3, performance results on Video-MME are relatively stable for both MLLMs. For this phenomenon, we assume that understanding fine spatial details is not vital for Video-MME. In contrast, the results on DREAM-1K increase a lot, which demonstrates the necessity of high resolution.
@@ 3.3 Intra-chunk Vision Encoder
where $n_{\text{V}}$ indicates the number of visual features per video chunk, and $d_{\text{V}}$ denotes the dimension of the visual features. We then adopt a standard ViT with parameter $\theta_{\text{ViT}}$ to capture high-level spatial-temporal features, denoted as ^$\hat{\mathcal{H}}_{i}$, within the $i$-th chunk. To manage the computational load and complexity for the downstream LLM module arising from a large number of tokens, we apply a 2x2 pooling layer on ^$\hat{\mathcal{H}}_{i}$ to obtain $\mathcal{H}_{i}\in\mathbb{R}^{n_{\text{V}}/4\times d_{\text{V}}}$.
@@ 3.4 Inter-chunk Feature Aggregator
[…]
@@ 3.5 Preprocessing
Video Preprocessing. The video processing strategy of Mavors varies based on the video length. Specifically, videos with short lengths are directly processed into chunks. To accommodate long videos, we employ an initial step of accelerated playback achieved through frame dropping, thereby reducing the total frame count to be compatible with Mavors processing limits. Specifically, the position IDs utilized by C-RoPE correspond to timestamps derived from the original, non-accelerated video timeline. This mechanism informs the model that the processed frames are not temporally contiguous. While alternative strategies for very long video comprehension exist, e.g., in-video Retrieval-Augmented Generation (RAG) [65], they represent an orthogonal direction to Mavors.
Meanwhile, Mavors could process videos with arbitrary resolutions and aspect ratios. Specifically, Mavors employs a dynamic resolution strategy to maintain the original aspect ratio of the video frames, avoiding distortion artifacts that can arise from fixed-shape resizing. The resized video frames roughly keep the original aspect ratio and match the number of pixels in the ViT’s pretraining images. For example, given the frames with the $(W_{\text{V}},H_{\text{V}})$ resolution and the ViT’s pretrained image resolution $(R_{v},R_{v})$, Mavors will rescale the frames into the resolution of $(R_{v}*\sqrt{W_{\text{V}}/H_{\text{V}}},R_{v}*\sqrt{H_{\text{V}}/W_{\text{V}}})$. We also resize the positional embedding of patches, following SigLIP [115]. Specifically, the positional embedding of the video chunk in the $(x,y)$ position, denoted as $E(x,y)$, will be formulated as:
@@
First, as image understanding tasks often require spatial details, we follow the image partition method in [110] and support dynamic resolution for processing high-resolution images, where the raw image will be partitioned into multiple sub-images and the size of these sub-images is supposed to match the number of pixels in the ViT’s pretraining. Specifically, we first determine the ideal number of sub-images $N_{s}=\left\lfloor(W_{\text{I}}\times H_{\text{I}})/R_{v}^{2}\right\rfloor$, where ($W_{\text{I}},H_{\text{I}}$) is the resolution of the original raw image and ($R_{v},R_{v}$) is the resolution of the ViT’s pretraining images. Next, we identify potential partition configurations by finding pairs of integers $(m,n)$, representing the number of columns and rows, respectively, such that their product equals the target number of slices $N_{s}$. These pairs form the set $\mathcal{C}_{N_{s}}=\{(m,n)|m\times n=N_{s},m,n\in\mathbb{Z}\}$. Then, we select the best configuration $(m^{*},n^{*})$ from $\tilde{C}=\mathcal{C}_{N_{s}-1}\cup\mathcal{C}_{N_{s}}\cup\mathcal{C}_{N_{s}+1}$ based on the following criteria:
@@ 4 Training Paradigm
Stage 1: Modality Alignment. As SigLIP’s training involved alignment with the T5 model [78], the first stage aims to align the semantic space of the vision encoder with the LLM’s semantic space. In this stage, we train the inter-chunk feature aggregator and the MLP projector, while keeping the LLM and the intra-chunk vision encoder frozen. Although the model exhibits only coarse video comprehension at this stage, the principal aim is to achieve modality alignment and instill basic temporal understanding. Therefore, we prioritize diverse, general-concept image-text pairs and short video-text pairs with low complexity (e.g., LAION [81] and PANDA-70M[12]), thereby avoiding excessively difficult data that could impede the development of foundational abilities.
@@
Stage 2: Multitask Instruction Tuning. In Stage 2, the primary objective is to adapt the model for a range of multi-modal tasks, leveraging data formats including text-only, single-image, multi-images, and complex video. Beyond standard CV tasks, we incorporate grounding tasks and temporal grounding tasks to enhance the model’s perception of spatio-temporal details. Similar to the practice in Qwen2.5VL [4], we find that representing bounding boxes using plain text coordinates yields performance comparable to using special tokens; consequently, we adopt the plain text representation. This stage also activates the sub-image partitioning paradigm to enhance the model’s image understanding capabilities. All model parameters are unfrozen and trained on a large dataset, allowing for extensive self-adjustment. Upon completion, the model possesses significant world knowledge, semantic understanding, and logical reasoning abilities, though its application is initially limited by the specific tasks and query formats encountered. Therefore, towards the end of this stage, we introduce more diverse data types, covering a broader spectrum of real-world task scenarios and textual query formulations.
Stage 3: DPO Training. Our empirical evaluations reveal that while the previously described training procedure yields strong leaderboard performance, the resulting model exhibits distinct patterns. Specifically, for QA tasks, the model tends to generate overly concise responses, likely due to extensive training on multiple-choice or short-answer datasets. Conversely, for descriptive tasks, the model fails to terminate generation appropriately. To mitigate these issues, we incorporate a Direct Preference Optimization (DPO) [77] stage following Stage 2. The preference dataset mainly covers three domains: open-ended QA, image captioning, and video captioning. More details can be found in Appendix A.
@@ 5.1 Experimental Setup
Implementation Details. The Mavors model utilizes Qwen2.5-7B as its language model module, with the intra-chunk vision encoder initialized using SigLIP weights. To balance effectiveness and efficiency, the frame count per video chunk, $F$, is set to 16. The inter-chunk feature aggregator consists of $L_{\text{Inter}}$=3 layers. The training is conducted on 416 A800-80GB GPUs. Given the model’s moderate size, we employed DeepSpeed with ZeRO stage 2 optimization. As mentioned in Section 4, the pre-training proceeded in three stages: Stage 1 used approximately 127 million samples with a global batch size of 6,656, taking 71 hours; Stage 1.5 used 52 million samples with a global batch size of 3,328, taking 177 hours; and Stage 2 used 19 million samples with a global batch size of 1,664, requiring 28 hours. The learning rates for the LLM and projector are set to 1e-5 in both Stage 1 and Stage 1.5, with a constant learning rate schedule applied during these phases. In Stage 2 and DPO, the learning rate was initialized at the same value (1e-5) as the preceding stages but followed a cosine decay schedule, gradually reducing to 1/10th of its initial value. Meanwhile, the learning rates for the inter-chunk feature aggregator and intra-chunk vision encoder remained fixed at 1/10th of the LLM’s learning rate across all training stages.
@@
For inference, Mavors is adapted using the vLLM framework [38]. Since Mavors requires comprehensive video encoding and frame preprocessing occurs on the CPU, the CPU processor can thus become a bottleneck. Recognizing that the intra-chunk vision encoder’s computation is a one-time GPU operation per video, with results stored in the LLM’s KV cache, we overlaps the pipeline. Specifically, the intra-chunk vision encoder and inter-chunk feature aggregator execute directly on the GPU, while the language model component leverages vLLM. This separation can effectively balance CPU-bound preprocessing, compute-intensive visual encoding (Intra/Inter), and language model inference. More details of the inference efficiency can be found in Appendix B.
@@
Baseline Models. We select several representative video models for performance comparison. We include GPT-4o-20240806 [32] and Gemini-1.5-Pro-002 [23] as the closed-source APIs baselines. Standard auto-regressive models using resolution-preserving frame sampling are represented by LLaVA-OneVision [43] and InternVL 2.5 [14]. For video understanding tasks, we add models based on: (a) high-performing sparse frame sampling (NVILA [61], LLaVA-Video [124]); (b) dense sampling with lower resolution (Qwen2.5-VL [4], Oryx-1.5 [60]); (c) dense sampling with token compression (VideoChat-Flash [47], VideoLLaMA3 [116]); and (d) slow-fast architecture, a special frame sampling strategy (Slow-fast MLLM [84]). Regarding image tasks, as some video-centric models either lack image input (e.g., VideoChat-Flash) or are not SOTA on image tasks, we include four strong models on QA/Caption benchmarks: GLM-4V [99], Qwen2.5-VL, DeepSeek-VL2 [105] and CogVLM2 [29]. Crucially, aside from prompt modifications, no benchmark-specific hyperparameters (e.g., frame sampling, resolution) were tuned during evaluation for any model, including Mavors.
@@
Benchmarks. Video understanding capabilities are assessed across general knowledge QA (MMWorld [28], PerceptionTest [74]), long-video QA (Video-MME [22], MLVU [126]), event understanding QA (MVBench [46], EventHallusion [117]), temporal understanding QA (TempCompass [58], VinoGround [118]), and captioning (DREAM-1K [96]). Image understanding evaluation includes comprehensive capabilities (MMMU [114]), cognitive understanding (MathVista [62], AI2D [37]), and captioning (CapsBench [52]). More experiment details can be found in Appendix C.

### v2

@@ 1 Introduction
Long-context video modeling stands as one of the most crucial capabilities within MLLMs [fuyu-8b, zhang2025videollama3, videochat-flash, Maaz2023VideoChatGPT]. This capability empowers MLLMs to proficiently manage hours-long movies, documentaries, and online video streams, all of which demand sophisticated long video processing. Recent advances in MLLMs perform well in short video understanding. However, it remains challenging to build MLLMs for processing extremely long videos (lasting for hours or even longer). The difficulty lies in how to enable MLLMs to efficiently understand the extremely long video context brought by long videos.
As shown in Figure 1, we have compared three mainstream types of video MLLMs with our method, and provided the video caption results of different methods for better illustration. Specifically, in Figure 1(a), these methods (e.g., LLaVA-Video [zhang2024video], InternVL 2.5 [chen2024expanding]) usually employ the sparse sampling strategy to decrease the number of frames and reduce the computation costs. However, these methods have a significant limitation, where many temporal contexts are lost as many frames are not sampled. Thus, the performance results of video-related tasks, which require detailed temporal contexts from many frames, are degraded a lot for these methods. When compared to methods in Figure 1(a), some methods (e.g., Oryx [liu2024oryx], Qwen2VL [Qwen2VL]) have introduced the strategy of dense sampling with low-resolution input in Figure 1(b). However, for these methods, many spatial contexts are lost as only the low-resolution frames are given, which also significantly degrade the results of video-related tasks requiring detailed spatial contexts, e.g., video captioning. Recently, in Figure 1(c), several works (e.g., VideoLLaMA 3 [zhang2025videollama3], VideoChat-Flash [videochat-flash]) have proposed token compression strategies (e.g., token merge or token dropping), which reduces tokens based on vector or pixel similarity and effectively preserves spatial-temporal features of large visual elements. However, token compression inevitably leads to the loss of information regarding small spatial objects, subtle temporal motions, and interactions among multiple objects, thereby posing challenges for understanding complex scenes.
@@ 2.1 MLLM Architecture
Current MLLMs employ two architectural strategies for visual processing. The first paradigm is based on cross-attention approach, which maintains frozen model parameters while establishing dynamic visual-language interactions through attention mechanisms [Alayrac2022FlamingoAV] . Alternatively, the second paradigm processes visual content through pretrained encoders (CLIP [radford2021clip], SigLIP [siglip]) before concatenating image tokens with text embeddings for unified language model processing [lin2023vila, liu2023llava, liu2023llava15, li2024llava, liu2024llavanext]. The second paradigm can be readily extensible to video analysis through sequential frame processing [2023videochat, zhang2025videollama3], and many architectural innovations for temporal modeling have been proposed [Jiang2025TokenEfficientLV, wang2025internvideo2, kangaroogroup].
@@ 2.2 MLLM for Video Understanding
Existing MLLMs have revealed divergent capabilities in temporal comprehension across different video durations. While existing systems demonstrate proficiency in minute-scale video analysis [2023videochat, videochat-flash, lin2023video], emerging efforts targeting hour-level sequences [gemini, wang2024longllava] face fundamental challenges. To address the challenges of long video modeling, current approaches primarily pursue two optimization directions: (1) context window expansion for large language models [gemini, wang2024longllava, Zhang2024LongCT, longvila] and (2) efficient token compression via spatial-temporal feature distillation [videoccam, li2024llamavid, shu2024video, song2023moviechat, TanKoala2024, Weng2024LongVLMEL]. For the first strategy, though theoretically enabling long-sequence processing, suffers from impractical computational overhead, which bring significant challenges for practical applications. In contrast, recent token compression methods like LLaMA-VID [li2024llamavid] achieve compression rates at the cost of discarding subtle details, which results in performance degradation on standard video understanding benchmarks. When compared to the existing works, our Mavors can directly process the raw videos to maintain spatial and temporal details well with acceptable computation costs.
@@ 3.1 Preliminaries
Necessity of Dense Sampling with High Resolution. As shown in Figure 2 and Figure 3, we have compared the results of two popular video MLLMs (i.e., Qwen2.5-VL-7B [qwen2.5vl] and Oryx-1.5-7B [liu2024oryx]) on two representative benchmarks (i.e., Video-MME [videomme] and DREAM-1K [wang2024tarsier]). Specifically, the Video-MME focuses on multiple-choice question answering based on video content and requires a better understanding of the temporal relations between different frames. DREAM-1K involves open-ended video captioning, where models must generate detailed descriptions of the main events in the video. Thus, both the spatial and temporal fine-grained details are important. In Figure 2, we observe that performance increases a lot when increasing the number of frames, which shows the necessity of dense sampling with more frames. In Figure 3, performance results on Video-MME are relatively stable for both MLLMs. For this phenomenon, we assume that understanding fine spatial details is not vital for Video-MME. In contrast, the results on DREAM-1K increase a lot, which demonstrates the necessity of high resolution.
@@ 3.3 Intra-chunk Vision Encoder
where $n_{\text{V}}$ indicates the number of visual features per video chunk, and $d_{\text{V}}$ denotes the dimension of the visual features. We then adopt a standard ViT with parameter $\theta_{\text{ViT}}$ to capture high-level spatial-temporal features, denoted as $\hat{\mathcal{H}}_{i}$, within the $i$-th chunk. To manage the computational load and complexity for the downstream LLM module arising from a large number of tokens, we apply a 2x2 pooling layer on $\hat{\mathcal{H}}_{i}$ to obtain $\mathcal{H}_{i}\in\mathbb{R}^{n_{\text{V}}/4\times d_{\text{V}}}$.
@@ 3.4 Inter-chunk Feature Aggregator
$\text{SDP}(q_{\iota}^{(j)},k_{\iota^{\prime}}^{(j)})$ $=\text{C-RoPE}(q_{\iota}^{(j)},k_{\iota^{\prime}}^{(j)};\lceil\frac{4\iota}{n_{\text{V}}}\rceil,\lceil\frac{4\iota^{\prime}}{n_{\text{V}}}\rceil)$ (4)
@@ 3.5 Preprocessing
Video Preprocessing. The video processing strategy of Mavors varies based on the video length. Specifically, videos with short lengths are directly processed into chunks. To accommodate long videos, we employ an initial step of accelerated playback achieved through frame dropping, thereby reducing the total frame count to be compatible with Mavors processing limits. Specifically, the position IDs utilized by C-RoPE correspond to timestamps derived from the original, non-accelerated video timeline. This mechanism informs the model that the processed frames are not temporally contiguous. While alternative strategies for very long video comprehension exist, e.g., in-video Retrieval-Augmented Generation (RAG) [luo2024video], they represent an orthogonal direction to Mavors.
Meanwhile, Mavors could process videos with arbitrary resolutions and aspect ratios. Specifically, Mavors employs a dynamic resolution strategy to maintain the original aspect ratio of the video frames, avoiding distortion artifacts that can arise from fixed-shape resizing. The resized video frames roughly keep the original aspect ratio and match the number of pixels in the ViT’s pretraining images. For example, given the frames with the $(W_{\text{V}},H_{\text{V}})$ resolution and the ViT’s pretrained image resolution $(R_{v},R_{v})$, Mavors will rescale the frames into the resolution of $(R_{v}*\sqrt{W_{\text{V}}/H_{\text{V}}},R_{v}*\sqrt{H_{\text{V}}/W_{\text{V}}})$. We also resize the positional embedding of patches, following SigLIP [siglip]. Specifically, the positional embedding of the video chunk in the $(x,y)$ position, denoted as $E(x,y)$, will be formulated as:
@@
First, as image understanding tasks often require spatial details, we follow the image partition method in [yao2024minicpmvgpt4vlevelmllm] and support dynamic resolution for processing high-resolution images, where the raw image will be partitioned into multiple sub-images and the size of these sub-images is supposed to match the number of pixels in the ViT’s pretraining. Specifically, we first determine the ideal number of sub-images $N_{s}=\left\lfloor(W_{\text{I}}\times H_{\text{I}})/R_{v}^{2}\right\rfloor$, where ($W_{\text{I}},H_{\text{I}}$) is the resolution of the original raw image and ($R_{v},R_{v}$) is the resolution of the ViT’s pretraining images. Next, we identify potential partition configurations by finding pairs of integers $(m,n)$, representing the number of columns and rows, respectively, such that their product equals the target number of slices $N_{s}$. These pairs form the set $\mathcal{C}_{N_{s}}=\{(m,n)|m\times n=N_{s},m,n\in\mathbb{Z}\}$. Then, we select the best configuration $(m^{*},n^{*})$ from $\tilde{C}=\mathcal{C}_{N_{s}-1}\cup\mathcal{C}_{N_{s}}\cup\mathcal{C}_{N_{s}+1}$ based on the following criteria:
@@ 4 Training Paradigm
Stage 1: Modality Alignment. As SigLIP’s training involved alignment with the T5 model [2020t5], the first stage aims to align the semantic space of the vision encoder with the LLM’s semantic space. In this stage, we train the inter-chunk feature aggregator and the MLP projector, while keeping the LLM and the intra-chunk vision encoder frozen. Although the model exhibits only coarse video comprehension at this stage, the principal aim is to achieve modality alignment and instill basic temporal understanding. Therefore, we prioritize diverse, general-concept image-text pairs and short video-text pairs with low complexity (e.g., LAION [laion400m] and PANDA-70M[chen2024panda]), thereby avoiding excessively difficult data that could impede the development of foundational abilities.
@@
Stage 2: Multitask Instruction Tuning. In Stage 2, the primary objective is to adapt the model for a range of multi-modal tasks, leveraging data formats including text-only, single-image, multi-images, and complex video. Beyond standard CV tasks, we incorporate grounding tasks and temporal grounding tasks to enhance the model’s perception of spatio-temporal details. Similar to the practice in Qwen2.5VL [qwen2.5vl], we find that representing bounding boxes using plain text coordinates yields performance comparable to using special tokens; consequently, we adopt the plain text representation. This stage also activates the sub-image partitioning paradigm to enhance the model’s image understanding capabilities. All model parameters are unfrozen and trained on a large dataset, allowing for extensive self-adjustment. Upon completion, the model possesses significant world knowledge, semantic understanding, and logical reasoning abilities, though its application is initially limited by the specific tasks and query formats encountered. Therefore, towards the end of this stage, we introduce more diverse data types, covering a broader spectrum of real-world task scenarios and textual query formulations.
Stage 3: DPO Training. Our empirical evaluations reveal that while the previously described training procedure yields strong leaderboard performance, the resulting model exhibits distinct patterns. Specifically, for QA tasks, the model tends to generate overly concise responses, likely due to extensive training on multiple-choice or short-answer datasets. Conversely, for descriptive tasks, the model fails to terminate generation appropriately. To mitigate these issues, we incorporate a Direct Preference Optimization (DPO) [Rafailov2023DirectPO] stage following Stage 2. The preference dataset mainly covers three domains: open-ended QA, image captioning, and video captioning. More details can be found in Appendix A.
@@ 5.1 Experimental Setup
Implementation Details. The Mavors model utilizes Qwen2.5-7B as its language model module, with the intra-chunk vision encoder initialized using SigLIP weights. To balance effectiveness and efficiency, the frame count per video chunk, $F$, is set to 16. The inter-chunk feature aggregator consists of $L_{\text{Inter}}$=3 layers. The training is conducted on 416 GPUs. Given the model’s moderate size, we employed DeepSpeed with ZeRO stage 2 optimization. As mentioned in Section 4, the pre-training proceeded in three stages: Stage 1 used approximately 127 million samples with a global batch size of 6,656, taking 71 hours; Stage 1.5 used 52 million samples with a global batch size of 3,328, taking 177 hours; and Stage 2 used 19 million samples with a global batch size of 1,664, requiring 28 hours. The learning rates for the LLM and projector are set to 1e-5 in both Stage 1 and Stage 1.5, with a constant learning rate schedule applied during these phases. In Stage 2 and DPO, the learning rate was initialized at the same value (1e-5) as the preceding stages but followed a cosine decay schedule, gradually reducing to 1/10th of its initial value. Meanwhile, the learning rates for the inter-chunk feature aggregator and intra-chunk vision encoder remained fixed at 1/10th of the LLM’s learning rate across all training stages.
@@
For inference, Mavors is adapted using the vLLM framework [kwon2023efficient]. Since Mavors requires comprehensive video encoding and frame preprocessing occurs on the CPU, the CPU processor can thus become a bottleneck. Recognizing that the intra-chunk vision encoder’s computation is a one-time GPU operation per video, with results stored in the LLM’s KV cache, we overlaps the pipeline. Specifically, the intra-chunk vision encoder and inter-chunk feature aggregator execute directly on the GPU, while the language model component leverages vLLM. This separation can effectively balance CPU-bound preprocessing, compute-intensive visual encoding (Intra/Inter), and language model inference. More details of the inference efficiency can be found in Appendix B.
@@
Baseline Models. We select several representative video models for performance comparison. We include GPT-4o-20240806 [hurst2024gpt] and Gemini-1.5-Pro-002 [gemini] as the closed-source APIs baselines. Standard auto-regressive models using resolution-preserving frame sampling are represented by LLaVA-OneVision [li2024llava] and InternVL 2.5 [chen2024expanding]. For video understanding tasks, we add models based on: (a) high-performing sparse frame sampling (NVILA [liu2024nvila], LLaVA-Video [zhang2024video]); (b) dense sampling with lower resolution (Qwen2.5-VL [qwen2.5vl], Oryx-1.5 [liu2024oryx]); (c) dense sampling with token compression (VideoChat-Flash [videochat-flash], VideoLLaMA3 [zhang2025videollama3]); and (d) slow-fast architecture, a special frame sampling strategy (Slow-fast MLLM [shi2025slow]). Regarding image tasks, as some video-centric models either lack image input (e.g., VideoChat-Flash) or are not SOTA on image tasks, we include four strong models on QA/Caption benchmarks: GLM-4V [wang2024cogvlm], Qwen2.5-VL, DeepSeek-VL2 [deepseek_vl2] and CogVLM2 [hong2024cogvlm2]. Crucially, aside from prompt modifications, no benchmark-specific hyperparameters (e.g., frame sampling, resolution) were tuned during evaluation for any model, including Mavors.
@@
Benchmarks. Video understanding capabilities are assessed across general knowledge QA (MMWorld [hemmworld], PerceptionTest [patraucean2023perception]), long-video QA (Video-MME [videomme], MLVU [MLVU]), event understanding QA (MVBench [li2024mvbench], EventHallusion [zhang2024eventhallusion]), temporal understanding QA (TempCompass [tempcompass], VinoGround [vinoground]), and captioning (DREAM-1K [wang2024tarsier]). Image understanding evaluation includes comprehensive capabilities (MMMU [yue2023mmmu]), cognitive understanding (MathVista [mathvista], AI2D [kembhavi2016diagram]), and captioning (CapsBench [liu2024playground]). More experiment details can be found in Appendix C.